OpenStudy (redneckoutlaw):

HELP Simplify: the square root of 2 (x − the square root of 5) Choose one answer. a. 2 times x times the square root of 2 b. 10 the square root of 2 times x c. the square root of 2 times x − the square root of 10 d. x times the square root of 2 − the square root of 10

OpenStudy (redneckoutlaw):

@Donblue22

OpenStudy (neer2890):

\[\sqrt{2}(x-\sqrt{5})=\sqrt{2}x-\sqrt{2}*\sqrt{5}=\sqrt{2}x-\sqrt{10}\] so d is correct.

OpenStudy (redneckoutlaw):

Thanks man, Can you help with some more

OpenStudy (neer2890):

you're welcome....:)

OpenStudy (redneckoutlaw):

Can you help with some more

OpenStudy (neer2890):

yeah.. ofcourse.., this site is for help only

OpenStudy (redneckoutlaw):

Im noot very good at math... Marks: 1 Find the value of 10 times the square root of 2 over the square root of 5.. Choose one answer. a. 2 b. 2 times the square root of 5. c. 2 the square root of 10. d. 2 over 5 times the square root of 10.

OpenStudy (neer2890):

\[\frac{ 10\sqrt{2} }{ \sqrt{5} }=\frac{ 2*5\sqrt{2} }{ \sqrt{5} }=\frac{ 2*\sqrt{5}*\sqrt{5}*\sqrt{2} }{ \sqrt{5} }= 2\sqrt{2*5}=2\sqrt{10}\]

OpenStudy (redneckoutlaw):

Do you simplifie that?

OpenStudy (neer2890):

i've already simplified it.

OpenStudy (redneckoutlaw):

But there is no answer with thaat answer??

OpenStudy (redneckoutlaw):

oh nvm

OpenStudy (redneckoutlaw):

Read it wrong Marks: 1 Find the value of The square root of the quantity 8 over 9. Choose one answer. a. 2 over 3. b. 1 and 1 third. c. 2 times the square root of 2, all over 3. d. 4 times the square root of 2, all over 3.

OpenStudy (neer2890):

\[\sqrt{\frac{ 8 }{ 9 }}=\sqrt{\frac{ 2*4 }{ 9 }}=\frac{ \sqrt{4}*\sqrt{2} }{ \sqrt{9} }=\frac{ 2\sqrt{2} }{ 3 }\]

OpenStudy (redneckoutlaw):

Marks: 1 Find the value of 3 times x over the square root of x cubed. Choose one answer. a. 3 over x. b. 3 times the square root of x. c. 3 over the square root of x. d. 3 times the square root of x, all over x.

OpenStudy (redneckoutlaw):

Im actuzally understanding theser. Thanks man

OpenStudy (neer2890):

you're welcome and try to do them first yourself. and show your work if you got stuck anywhere.

OpenStudy (neer2890):

\[\frac{ 3x }{ \sqrt{x ^{3}} }=\frac{ 3x }{ x ^{\frac{ 3 }{ 2 }} }=3x*x ^{-\frac{ 3 }{ 2 }}=3x ^{(1-\frac{ 3 }{ 2 })}\]

OpenStudy (neer2890):

\[=3x ^{\frac{ -1 }{ 2 }}=\frac{ 3 }{ \sqrt{x} }\]

OpenStudy (redneckoutlaw):

Marks: 1 Find the value of 3 over the square root of 3.. Choose one answer. a. 3 b. 3 times the square root of 3. c. the square root of 3. d. the square root of 3 over 3.

OpenStudy (redneckoutlaw):

@Donblue22

OpenStudy (redneckoutlaw):



OpenStudy (donblue22):

c

OpenStudy (redneckoutlaw):

You positive? ind the value of 2 times the square root of 3, all over the square root of 10. in simplest form. Choose one answer. a. 2 times the square root of 30. b. 2 times the square root of 3, all over 5. c. 2 times the square root of 30, all over 10. d. the square root of 30, all over 5.

OpenStudy (donblue22):

ya remmeber for that kind of problem mulitply the top and bottom by the redical on the botttom

OpenStudy (donblue22):

pretty sure this on is also c

OpenStudy (redneckoutlaw):

Find the value of the square root of 6 over 5 times the square root of 3. in simplest form. Choose one answer. a. 2 over 5. b. 3 over 4. c. the square root of 15 over 5. d. the square root of 2 over 5.

OpenStudy (redneckoutlaw):

@IMStuck

OpenStudy (imstuck):

\[\frac{ \sqrt{6} }{ 5 }*\sqrt{3}\]Is that it?

OpenStudy (redneckoutlaw):